OpenStudy (anonymous):

The combination of gaseous atoms of non-metal X and of hydrogen atom to form gaseous X2H6 molecules. 2X(g) + 6H (g). ----> X2H6 (g) The bond energy of an X-H is 395kJ/mil. What's the bond energy of X-X?

OpenStudy (chmvijay):

did they have mentioned any other data please check once again

OpenStudy (chmvijay):

you have to tell me the delta H for the following reaction it is missing in the question

OpenStudy (chmvijay):

really some value is missing in this :)

OpenStudy (anonymous):

Oh, yes! I missed that. That's -2775KJ/mol

OpenStudy (chmvijay):

ok ok

OpenStudy (chmvijay):

in this X2 H6 you have 6 X-H bonds and 1 X-X bond right

OpenStudy (anonymous):

Yeah I know that already

OpenStudy (chmvijay):

energy liberated bcoz of both 6-X-H bond +1 X-X bond is -2775 KJ but if u calculate for the 6-X-H bonds it is =6*395 = 2370 KJ therefore remaining willbe liberated when X-X bonds are formed =2775-2370 = 405 KJ

OpenStudy (chmvijay):

did u get it :)

OpenStudy (anonymous):

Sign changing? How?

OpenStudy (chmvijay):

whenever bonds are formed energy will be released hence- ve sign :)

OpenStudy (chmvijay):

6 X-H + X-X = -2775 KJ -2370 + X-X = -2775 KJ X-X =-2775+2370 = -405 KJ

OpenStudy (chmvijay):

bond energy is positive (+) or may be negative(-) if bonds are formed u use -ve sign bcoz energy is liberated in this phenomenon if bonds has to be broken out then use a +ve sign bcoz energy is needed to break it

OpenStudy (anonymous):

I got it now. Thank you so much. :)